kekeman:

Expand the following using the Binomial Theorem and Pascal's triangle: (x + 2)^6

minustempo:

Do you know the binomial theorem? It's (x+y)^z = C(z,0)*x^z*y^0 + C(z,1)*x^(z-1)*y^1 + C(z,2)*x^(z-2)*y^2 ... + C(z,z-1)*x^1*y^(z-1) + C(z,z)*x^0*y^z

minustempo:

C(n,r) is the number of combinations for choosing r things from n things

minustempo:

Pascal's triangle is the 1 1 1 1 2 1 1 3 3 1 that thing

minustempo:

Alright it's x^6 + 12x^5 + 60x^4 + 160x^5 + 240x^2 + 192x + 128

kekeman:

Yes i know i am going to show you my steps

kekeman:

= x^6 2^0 + 6(x^5)(2)^1 + 15(x^4)(2^2) + 20(x^3)(2^3) + 15(x^2)(2^4) + 6(x)(2^5) + 1(2^6) = x^6 + 12x^5 + 60x^4 + 160x^3 + 240x^2 + 192x + 64

minustempo:

Yeah that's right

kekeman:

Ok thanks